OpenStudy (anonymous):

Tickets to the museum cost $18 for adults and $12 for children. A group of 20 people went to the museum, and the tickets cost $324. The system of equations models this situation, where x is the number of adults and y is the number of children. How many adults and how many children were in the group? A. 8 adults and 12 children B. 10 adults and 10 children C. 12 adults and 8 children D. 14 adults and 6 children

OpenStudy (anonymous):

system of equation is x+y=20 18x+12y=324

OpenStudy (anonymous):

@sissyedgar

OpenStudy (anonymous):

ANSWER

OpenStudy (anonymous):

HELLO

OpenStudy (4everaddicted2anime):

Isolate the x in the top equation

OpenStudy (anonymous):

I DONT GET THIS AT ALLL

OpenStudy (sissyedgar):

The answer would be D... 14 (adults) multiplied by 18 (price for adults) equals 252! 6 (children) multiplied by 12 (price for children) equals 72! 252 + 72 = 324 ღ

OpenStudy (4everaddicted2anime):

Okay. The top equation is the simplest so it would be best to work with that first. The equation is x+y=20. To isolate the x you need to move everything from the left to the right, not including x. To do this you must subtract y from both sides to get x=20-y. From here you plug that into the second equation to get your y coordinate